OpenStudy (tiff9702):

Evaluate the expression when x = 8 and y = 1. x/2y A. 1/8 B. 8 C. 1/4 D. 4

OpenStudy (tiff9702):

@.Sam.

OpenStudy (arindameducationusc):

8/2

OpenStudy (anonymous):

you have to substitute x=8 and y =1

OpenStudy (arindameducationusc):

and 8/2 is ?

OpenStudy (arindameducationusc):

agreed with @GIL.ojei

OpenStudy (arindameducationusc):

4

OpenStudy (tiff9702):

I know I was Never good at this

OpenStudy (anonymous):

that will give\[\frac{ x }{ 2y } =\frac{ 8 }{ 2(1) }=\frac{ 8 }{ 2 }=4\]